# SAKE: Towards Editing Auditory Attribute Knowledge of Large Audio-Language Models

### v1

@@ Appendix D More Details and Evaluation Prompts for LLM-as-a-judge
FT (LLM) ALL 100.00 99.94 100.00 100.00 99.83 67.42 91.67 10.44 83.33 70.00 74.58 24.67
Animal 100.00 99.89 100.00 100.00 99.67 64.00 93.00 8.67 86.33 68.00 70.00 22.67
Emotion 100.00 100.00 100.00 100.00 100.00 59.00 93.67 8.00 66.00 68.33 80.00 20.67
Gender 100.00 100.00 100.00 100.00 100.00 83.78 85.67 - 92.33 73.33 77.67 26.33
Language 100.00 99.89 100.00 100.00 99.67 67.00 94.33 14.67 88.67 70.33 70.67 29.00
FT (Audio) ALL 100.00 81.86 99.00 77.17 69.42 90.53 96.83 80.11 90.00 92.58 100.00 50.67
Animal 100.00 84.44 99.00 81.00 73.33 94.25 98.00 94.00 93.33 91.67 100.00 48.67
Emotion 100.00 71.67 99.00 62.67 53.33 85.08 98.33 67.00 84.33 90.67 100.00 43.00
Gender 100.00 97.44 100.00 96.00 96.33 94.56 98.00 - 92.00 93.67 100.00 63.00
Language 100.00 73.89 98.00 69.00 54.67 89.25 93.00 79.33 90.33 94.33 100.00 48.00
KE ALL 95.50 86.67 92.00 87.67 80.33 83.47 89.83 61.44 87.25 89.83 84.92 27.58
Animal 97.00 88.22 89.33 94.00 81.33 88.42 93.67 76.33 92.33 91.33 81.33 24.67
Emotion 98.33 76.67 91.33 74.33 64.33 78.00 97.33 52.67 72.00 90.00 87.00 23.33
Gender 87.67 89.22 90.33 88.33 89.00 84.78 72.67 - 93.67 88.00 85.00 39.00
Language 99.00 92.56 97.00 94.00 86.67 83.00 95.67 55.33 91.00 90.00 86.33 23.33
MEND ALL 100.00 95.33 98.92 95.92 91.17 83.27 98.50 49.33 85.42 91.33 86.75 27.17
Animal 100.00 97.33 98.00 98.33 95.67 84.50 97.00 64.00 89.67 87.33 83.67 22.67
Emotion 100.00 89.89 99.00 92.00 78.67 75.67 100.00 42.00 70.33 90.33 89.00 23.67
Gender 100.00 95.89 100.00 94.33 93.33 94.89 99.00 - 92.33 93.33 88.67 37.00
Language 100.00 98.22 98.67 99.00 97.00 80.92 98.00 42.00 89.33 94.33 85.67 25.33
UnKE ALL 98.60 98.53 99.08 98.83 97.67 67.49 91.42 12.34 82.92 69.50 71.58 30.18
Animal 98.00 97.67 97.67 96.33 99.00 63.50 88.00 10.67 86.67 68.67 73.30 21.70
Emotion 96.70 97.33 99.00 99.67 93.33 62.25 95.33 18.67 64.67 70.33 70.30 33.30
Gender 100.00 99.56 100.00 99.67 99.00 82.67 90.00 - 91.67 66.33 72.00 36.70
Language 99.70 99.56 99.67 99.67 99.33 65.33 92.33 7.67 88.67 72.67 70.70 29.00
I-IKE ALL 10.33 7.11 10.33 5.67 5.33 87.51 94.75 89.33 93.00 73.42 55.00 28.92
Animal 6.67 7.11 7.00 7.33 7.00 90.08 96.33 94.00 94.33 75.67 54.33 25.00
Emotion 19.67 13.22 19.67 9.00 11.00 82.58 97.00 77.67 83.33 72.33 58.67 23.00
Gender 9.67 3.11 8.67 0.67 0.00 88.22 94.00 - 99.00 71.67 56.00 44.00
Language 5.33 5.00 6.00 5.67 3.33 89.33 91.67 96.33 95.33 74.00 51.00 23.67
IE-IKE ALL 8.00 6.58 8.50 5.75 5.50 82.89 91.25 86.56 89.67 65.00 50.00 27.50
Animal 6.33 6.67 7.33 6.00 6.67 84.17 90.33 90.67 88.33 67.33 51.33 26.00
Emotion 9.67 11.67 11.67 11.00 12.33 79.50 96.67 76.33 80.33 64.67 51.00 21.67
Gender 11.00 3.78 10.67 0.67 0.00 83.67 89.00 - 99.00 63.00 50.67 40.33
Language 5.00 4.22 4.33 5.33 3.00 84.42 89.00 92.67 91.00 65.00 47.00 22.00
Appendix E Detailed Results for Each Auditory Attribute under Single Editing
@@
We compare the performance for each auditory attribute (animal sound, speaker emotion, speaker gender, and spoken language) under single editing here, as shown in Table 6 and Table 7 for DeSTA2.5-Audio and Qwen2-Audio, respectively.
@@
Reliability. All methods except I-IKE and IE-IKE exhibit consistently high reliability across all attributes, suggesting that they can effectively update the edited knowledge on both LALMs. In contrast, the performance of I-IKE and IE-IKE varies by attribute and model, which we attribute to differences in the original model’s ability to perceive auditory attributes.
@@
Generality. For most methods, their generality remains relatively stable across different auditory attributes. However, FT (Audio) exhibits a notable exception: although it achieves consistently high reliability, its generality on emotion attributes drops the most compared to its reliability on both LALMs. This suggests that editing the modality connector makes it harder for the model to extend edited knowledge to similar inputs within emotion.
@@
Locality. Most methods achieve relatively higher average performance on the gender attribute, owing to the inapplicability of type 2 audio locality, which was identified in Sec. 5.1 as the most difficult to preserve, thereby inflating the overall average. For attributes other than gender, most methods perform comparably across attributes on Qwen2-Audio. In contrast, on DeSTA2.5-Audio, although performance on type 1 and type 4 is generally similar, higher audio locality is observed for language in type 2 and type 3. This indicates that unrelated auditory knowledge regarding spoken language is less susceptible to disruption when edits are applied on DeSTA2.5-Audio. Regarding text locality, most methods demonstrate comparable performance across attributes on both LALMs.
@@
Portability. Different methods result in varying portability performance across attributes on the two LALMs. A consistent observation on both models is that FT (Audio) achieves higher portability scores for all atributes, especially gender, suggesting that editing the modality connector may more effectively propagate the edited knowledge to other interconnected knowledge, particularly that related to speaker gender.
Appendix F Detailed Results under Sequential Editing
@@
We provide the statistics of the sequential editing results in Table 8 and 9, with the corresponding line charts shown in Figure 3(a) and Figure 3(b) for DeSTA2.5-Audio and Qwen2-Audio, respectively.
@@
Table 8: Original result of the four metrics of different editing methods on DeSTA2.5-Audio under sequential editing. For generality and audio locality, we present the averaged results. (%)
@@
3 32.00 34.00 54.92 54.00 34.00
4 32.00 39.33 53.37 54.00 38.00
5 26.00 32.00 52.33 54.00 34.00
I-IKE 0 60.00 54.67 65.28 50.00 56.00
1 56.00 50.67 60.10 52.00 52.00
2 56.00 54.67 61.66 68.00 54.00
3 56.00 52.00 62.69 52.00 48.00
4 62.00 56.67 61.14 52.00 46.00
5 62.00 56.67 59.07 56.00 54.00
Table 9: Original result of the four metrics of different editing methods on Qwen2-Audio under sequential editing. For generality and audio locality, we present the averaged results. (%)

### v2

@@ Appendix F Detailed Results for Each Auditory Attribute under Single Editing
FT (LLM) ALL 100.00 99.94 100.00 100.00 99.83 67.40 91.50 10.33 83.33 70.17 75.58 24.00
Animal 100.00 99.89 100.00 100.00 99.67 64.17 93.00 8.33 86.33 69.00 70.33 22.33
Emotion 100.00 100.00 100.00 100.00 100.00 58.83 93.33 8.00 66.00 68.00 80.33 19.00
Gender 100.00 100.00 100.00 100.00 100.00 83.67 85.33 - 92.33 73.33 79.67 26.00
Language 100.00 99.89 100.00 100.00 99.67 67.00 94.33 14.67 88.67 70.33 72.00 28.67
FT (Audio) ALL 100.00 81.81 99.00 77.00 69.42 90.49 96.75 80.00 90.00 92.58 100.00 49.92
Animal 100.00 84.22 99.00 80.33 73.33 94.17 98.00 93.67 93.33 91.67 100.00 48.33
Emotion 100.00 71.67 99.00 62.67 53.33 85.00 98.33 67.00 84.33 90.33 100.00 42.33
Gender 100.00 97.44 100.00 96.00 96.33 94.67 98.00 - 92.00 94.00 100.00 62.33
Language 100.00 73.89 98.00 69.00 54.67 89.17 92.67 79.33 90.33 94.33 100.00 46.67
KE ALL 95.33 86.69 92.00 87.92 80.17 83.36 89.83 61.44 87.25 89.42 85.67 26.75
Animal 97.00 88.11 89.33 94.00 81.00 88.33 93.67 76.33 92.33 91.00 82.67 23.00
Emotion 98.33 76.67 91.33 74.33 64.33 77.92 97.33 52.67 72.00 89.67 86.67 22.67
Gender 87.00 89.44 90.33 89.33 88.67 84.44 72.67 - 93.67 87.00 85.67 38.33
Language 99.00 92.56 97.00 94.00 86.67 83.00 95.67 55.33 91.00 90.00 87.67 23.00
MEND ALL 100.00 95.31 98.92 95.92 91.08 83.29 98.58 49.44 85.42 91.25 87.25 26.25
Animal 100.00 97.22 98.00 98.33 95.33 84.50 97.00 64.33 89.67 87.00 84.67 21.33
Emotion 100.00 89.89 99.00 92.00 78.67 75.67 100.00 42.00 70.33 90.33 89.67 23.33
Gender 100.00 95.89 100.00 94.33 93.33 94.89 99.00 - 92.33 93.33 89.33 36.00
Language 100.00 98.22 98.67 99.00 97.00 81.00 98.33 42.00 89.33 94.33 85.33 24.33
UnKE ALL 98.67 98.64 99.42 98.75 97.75 68.00 92.08 12.22 83.00 67.17 72.84 26.58
Animal 97.00 97.89 98.67 95.67 99.33 63.42 91.00 9.33 87.00 66.33 72.67 20.33
Emotion 98.33 97.33 99.00 99.67 93.33 62.00 95.33 18.67 64.67 69.33 71.00 16.33
Gender 100.00 99.56 100.00 99.67 99.00 82.33 90.00 - 91.67 65.33 75.67 38.00
Language 99.33 99.78 100.00 100.00 99.33 64.25 92.00 8.67 88.67 67.67 72.00 31.67
I-IKE ALL 10.33 7.00 10.25 5.58 5.17 87.47 94.75 89.56 93.00 73.08 55.67 28.50
Animal 7.00 6.89 6.67 7.33 6.67 89.75 96.33 94.33 94.33 74.00 57.33 24.00
Emotion 19.33 13.11 19.67 9.00 10.67 82.75 97.00 78.00 83.33 72.67 58.00 22.00
Gender 9.67 3.00 8.67 0.33 0.00 88.67 94.33 - 99.00 72.67 57.33 45.00
Language 5.33 5.00 6.00 5.67 3.33 89.00 91.33 96.33 95.33 73.00 50.00 23.00
IE-IKE ALL 7.83 6.44 8.33 5.67 5.33 82.82 91.25 86.67 89.67 64.67 51.17 27.58
Animal 6.33 6.22 6.67 6.00 6.00 84.17 90.33 91.00 88.33 67.00 52.67 26.33
Emotion 9.00 11.56 11.67 10.67 12.33 79.50 96.67 76.33 80.33 64.67 54.33 21.67
Gender 11.00 3.78 10.67 0.67 0.00 83.44 89.00 - 99.00 62.33 52.00 40.00
Language 5.00 4.22 4.33 5.33 3.00 84.33 89.00 92.67 91.00 64.67 45.67 22.33
WISE ALL 100.00 99.34 99.92 99.17 98.92 70.04 92.67 6.22 82.84 77.75 82.09 27.84
Animal 100.00 99.67 99.67 99.67 99.67 64.00 90.67 3.67 86.00 75.67 79.67 6.67
Emotion 100.00 98.78 100.00 98.67 97.67 60.58 93.33 9.67 65.00 74.33 82.00 21.67
Gender 100.00 98.89 100.00 98.33 98.33 88.89 94.00 - 91.67 81.00 83.67 41.00
Language 100.00 100.00 100.00 100.00 100.00 66.67 92.67 5.33 88.67 80.00 83.00 42.00
Table 10: Detailed results of the four metrics of each auditory attribute across different editing methods on AF under single editing. Attr. denotes auditory attributes, and Port. denotes portability. For generality and audio locality, Avg. indicates the average performance across all types of the corresponding metric. (%)
Avg. Type 1 Type 2 Type 3 Avg. Type 1 Type 2 Type 3 Type 4
FT (LLM) ALL 100.00 93.31 96.67 94.17 89.08 87.67 98.00 69.89 85.42 92.92 97.00 31.67
Animal 100.00 91.67 92.33 97.33 85.33 89.83 98.33 79.33 89.67 92.00 95.67 29.00
Emotion 100.00 88.89 98.33 85.00 83.33 78.08 98.00 53.67 69.67 91.00 98.00 17.33
Gender 100.00 98.89 100.00 98.33 98.33 93.89 97.33 - 88.67 95.67 98.33 53.00
Language 100.00 93.78 96.00 96.00 89.33 90.42 98.33 76.67 93.67 93.00 96.00 27.33
FT (Audio) ALL 100.00 76.08 96.67 74.00 57.58 90.24 96.92 78.56 89.33 93.25 100.00 49.50
Animal 100.00 59.33 91.00 55.33 31.67 96.08 98.67 94.67 97.33 93.67 100.00 46.33
Emotion 100.00 63.00 99.33 49.33 40.33 89.25 98.67 82.33 83.00 93.00 100.00 38.00
Gender 100.00 91.44 99.67 96.67 78.00 92.89 96.67 - 87.67 94.33 100.00 65.67
Language 100.00 90.56 96.67 94.67 80.33 83.42 93.67 58.67 89.33 92.00 100.00 48.00
KE ALL 100.00 98.89 99.92 98.75 98.00 81.38 91.50 55.78 78.50 93.33 95.83 31.92
Animal 100.00 99.22 99.67 99.67 98.33 85.92 92.00 70.00 90.00 91.67 93.33 28.33
Emotion 100.00 97.11 100.00 96.33 95.00 71.75 90.00 36.00 68.00 93.00 97.00 19.00
Gender 100.00 100.00 100.00 100.00 100.00 84.33 95.67 - 62.67 94.67 97.33 52.33
Language 100.00 99.22 100.00 99.00 98.67 84.25 88.33 61.33 93.33 94.00 95.67 28.00
MEND ALL 99.92 93.92 99.08 92.42 90.25 90.64 99.25 79.22 85.58 95.67 98.67 32.08
Animal 100.00 97.67 98.33 99.33 95.33 91.58 99.67 81.33 90.00 95.33 98.67 30.67
Emotion 100.00 87.44 99.00 82.00 81.33 84.42 99.00 70.67 72.67 95.33 99.00 17.00
Gender 100.00 93.56 100.00 90.33 90.33 93.44 99.00 - 85.67 95.67 99.33 53.67
Language 99.67 97.00 99.00 98.00 94.00 93.83 99.33 85.67 94.00 96.33 97.67 27.00
UnKE ALL 99.92 91.17 98.09 99.25 94.17 76.03 93.33 31.22 83.33 82.22 81.17 31.75
Animal 99.67 95.67 94.67 100.00 92.33 75.73 91.33 41.00 88.67 81.94 77.67 27.67
Emotion 100.00 95.33 99.67 97.00 89.33 67.00 93.67 27.67 67.33 79.33 80.67 17.33
Gender 100.00 99.67 100.00 100.00 99.00 87.21 94.00 - 84.33 83.28 85.67 54.00
Language 100.00 98.00 98.00 100.00 96.00 74.17 94.33 25.00 93.00 84.33 80.67 28.00
I-IKE ALL 25.00 15.08 25.17 9.50 10.58 87.07 95.83 88.00 89.17 75.50 81.17 37.08
Animal 12.67 12.11 11.33 10.67 14.33 89.25 94.67 89.00 93.33 80.00 81.67 37.67
Emotion 24.00 9.11 24.67 1.33 1.33 83.25 97.00 83.33 80.00 72.67 80.67 22.00
Gender 55.33 32.67 57.33 20.00 20.67 86.56 96.00 - 89.00 74.67 81.00 59.67
Language 8.00 6.44 7.33 6.00 6.00 89.08 95.67 91.67 94.33 74.67 81.33 29.00
IE-IKE ALL 58.83 58.36 58.25 58.25 58.58 40.42 34.50 19.11 60.92 41.83 76.67 40.67
Animal 38.33 41.11 42.00 38.33 43.00 39.67 37.33 20.00 53.33 48.00 78.00 41.00
Emotion 51.33 49.33 50.00 49.33 48.67 35.58 34.33 21.33 48.00 38.67 77.67 25.67
Gender 99.33 98.56 99.33 98.33 98.00 51.78 29.33 - 84.33 41.67 75.67 61.33
Language 46.33 44.44 41.67 47.00 44.67 37.50 37.00 16.00 58.00 39.00 75.33 34.67
WISE ALL 100.00 77.69 85.75 83.67 63.67 91.18 97.92 84.78 87.75 92.07 96.08 30.75
Animal 100.00 83.00 80.00 95.67 73.33 92.99 97.67 89.67 90.67 93.98 97.00 30.00
Emotion 100.00 79.33 96.00 74.00 68.00 84.17 99.00 74.67 73.33 89.67 96.00 18.00
Gender 100.00 76.00 97.33 74.33 56.33 93.55 97.00 - 91.33 92.31 97.00 51.00
Language 100.00 72.44 69.67 90.67 57.00 94.00 98.00 90.00 95.67 92.33 94.33 24.00
@@
We compare the performance for each auditory attribute (animal sound, speaker emotion, speaker gender, and spoken language) under single editing here, as shown in Table 8, Table 9, and Table 10 for DeSTA, Qwen, and AF, respectively.
@@
Reliability remains high across all attributes. All methods except I-IKE and IE-IKE exhibit consistently high reliability across all attributes, suggesting that they can effectively update the edited knowledge on these LALMs. In contrast, the performance of I-IKE and IE-IKE varies by attributes and models, which we attribute to differences in the models’ in-context learning ability.
@@
Editing the modality connector limits emotional generalization. For most methods, the generality remains relatively stable across different auditory attributes. However, FT (Audio) exhibits a notable exception: while it consistently achieves high reliability, its generality on emotion attributes degrades the most compared to its reliability on DeSTA (99.67% $\rightarrow$ 78.56%) and Qwen (100.00% $\rightarrow$ 71.67%), and the second most on AF (100.00% $\rightarrow$ 63.00%), among all attributes. This suggests that editing the modality connector makes it harder for the model to extend edited knowledge to similar inputs within emotion.
@@
Audio locality varies by attributes and models. Most methods achieve relatively higher average performance on the gender attribute, owing to the inapplicability of type 2 audio locality, which was identified in Sec. 5.2 as the most difficult to preserve, thereby inflating the overall average. For attributes other than gender, most methods perform comparably across attributes on Qwen. In contrast, on DeSTA, higher audio locality is observed for language in types 2 and 3, while on AF, lower audio locality is observed for emotion in type 2 across most methods. These results indicate attribute-dependent differences in preservation difficulty across LALMs. Regarding text locality, most methods demonstrate comparable performance across attributes on these LALMs.
@@
FT (Audio) achieves consistently high portability across attributes. Different methods result in varying portability performance across attributes on these LALMs. A consistent observation on these models is that FT (Audio) achieves higher portability scores for all attributes, suggesting that editing the modality connector may more effectively propagate the edited knowledge to other interconnected knowledge.
Appendix G Detailed Results under Sequential Editing
@@
We provide the statistics of the sequential editing results in Table 11, Table 12, and Table 13, with the corresponding line charts shown in Figure 4(a), Figure 4(b), and Figure 4(c) for DeSTA, Qwen, and AF, respectively.
@@
Table 11: Original result of the four metrics of different editing methods on DeSTA under sequential editing. For generality and audio locality, we present the averaged results. (%)
@@
3 32.00 34.00 54.92 58.00 34.00
4 32.00 39.33 52.85 58.00 38.00
5 26.00 32.00 52.85 56.00 34.00
I-IKE 0 58.00 54.67 65.80 52.00 58.00
1 56.00 50.67 60.62 54.00 52.00
2 56.00 54.67 61.14 66.00 54.00
3 56.00 52.00 62.18 54.00 46.00
4 62.00 56.67 60.10 54.00 46.00
5 62.00 56.67 59.07 56.00 52.00
WISE 0 100.00 100.00 41.45 88.00 20.00
1 78.00 75.33 40.41 94.00 16.00
2 76.00 68.00 37.82 98.00 18.00
3 74.00 66.00 38.34 96.00 14.00
4 70.00 61.33 37.82 90.00 16.00
5 62.00 54.67 37.82 92.00 16.00
Table 12: Original result of the four metrics of different editing methods on Qwen under sequential editing. For generality and audio locality, we present the averaged results. (%)